OpenStudy (anonymous):

Graph the quadratic functions y = -2x^2 and y = -2x^2 + 4.. after that compare and contrast the shape and position of both graphs. Help please

OpenStudy (badhi):

If the quadratic function is,\[f(x) = ax^2+bx+c\], adjust it like the following, \(f(x)=a\left(x^2+\frac ba x + \frac c a \right)\) \(=a\left[\left(x+\frac{b}{2a}\right)^2+\left(\frac c a-\frac {b^2}{4a^2}\right)\right] ---(1)\) by looking at the values of \(a,b,c\) we can predict the graph. If \(a>0\) the graph will have a single minima else a single maxima the minimum point is when \(x=\frac{-b}{2a}\) To find the x values where the graph crosses the x axis find the roots of \(f(x)=0\)

OpenStudy (anonymous):

i don't really understand..

OpenStudy (anonymous):

can you please help me @iambatman